OpenStudy (anonymous):

PLEASE HELP! Which number is equal to 7(3.5 x 10^4) Written in scientific notation? 24.5 x 70^4 2.45 x 10^5 24.5 x 10^4 1.05 x 10^5

OpenStudy (anonymous):

7(3.5*10^4)) =7*3.5*10^4 =24.5*10^4 =2.45*10^5

OpenStudy (anonymous):

Thanks :) It was right .

OpenStudy (anonymous):

yw